# MultiOCR-QA: Dataset for Evaluating Robustness of LLMs in Question Answering on Multilingual OCR Texts

### v1

@@ Abstract.
Optical Character Recognition (OCR) plays a crucial role in digitizing historical and multilingual documents, yet OCR errors -imperfect extraction of the text, including character insertion, deletion and permutation- can significantly impact downstream tasks like question-answering (QA). In this work, we introduce a multilingual QA dataset MultiOCR-QA, designed to analyze the effects of OCR noise on QA systems’ performance. The MultiOCR-QA dataset comprises 60K question-answer pairs covering three languages, English, French, and German. The dataset is curated from OCR-ed old documents, allowing for the evaluation of OCR-induced challenges on question answering. We evaluate MultiOCR-QA on various levels and types of OCR errors to access the robustness of LLMs in handling real-world digitization errors. Our findings show that QA systems are highly prone to OCR induced errors and exhibit performance degradation on noisy OCR text.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 13–18, 2025; Padova, ITisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Question answeringccs: Information systems Content analysis and feature selection
@@ 1. Introduction
Optical Character Recognition (OCR) technology has played a crucial role in preserving and providing access to historical texts. Over the past decade, significant advancements in OCR have improved text recognition accuracy, leading to the development of large-scale digital libraries of historical texts (TERRAS, 2011). These libraries serve as valuable resources for researchers, historians, and the general public, granting access to old manuscripts, newspapers, and other archival materials. Historical documents hold a wealth of knowledge, offering insights into past events, cultures, and people. Many professionals such as historians, journalists, or sociologists rely on these collections for various research and analysis tasks.
Historical corpora have been widely used for numerous natural language processing (NLP) tasks, including Named Entity Recognition (NER), topic modeling, text classification, neural ranking, and understanding semantic changes in language over time (Ehrmann et al., 2020, 2023; Yang et al., 2011; Liebeskind and Liebeskind, 2020; Subramani et al., 2020). Among these, question answering (QA) provides a direct approach to extracting precise and relevant information from historical texts. Unlike broader text-processing tasks, QA focuses on retrieving specific answers to user queries, making it an essential tool for efficiently accessing historical knowledge.
@@
Despite the advancements in OCR technology, several challenges persist. OCR-generated text, referred as RawOCR text in this paper, often contains errors due to the degraded conditions of historical documents. Factors like faded ink and paper, irregular fonts, physical damage, and printing inconsistencies cause recognition errors that negatively impact downstream NLP tasks such as information retrieval, machine translation, and QA systems. Since QA models heavily depend on the quality of the input text, errors in raw OCR text can significantly affect the accuracy and reliability of generated answers.
@@
For instance, in German, a passage with OCR error ”Der Bericht der Lagsatzungsgesandtschaft wird verlesen undvon Hrn. Bürgermeistet Mousson als erstem Gesandten desStandes Zürich mit einigen Bemerkungen begleitet. [^1] English Translation: The report of the legislative mission is read out and accompanied by a few comments from Mr. Mayor Mousson, the first envoy of the Zurich state.” contains multiple insertion, deletion, and substitution errors, such as ”Lagsatzungsgesandtschaft” (should be ”Tagsatzungsgesandtschaft”) and ”Bürgermeistet” (should be ”Bürgermeister”). When a QA model encounters such a noisy OCR text, it may generate an incorrect answer. For example, given the question: ”Wer hat den Bericht der Tagsatzungsgesandtschaft verlesen?” (Who read the report of the parliamentary delegation?) the model incorrectly responds ”Der Bericht der Tagsatzungsgesandtschaft wurde von Hrn. Bürgermeistet Mousson verlesen.” (The report of the parliamentary delegation was read out by Mr. Mayor Mousson.) which retains the OCR error and potentially misleads the QA system. This example highlights how even minor OCR errors can significantly affect QA accuracy, leading to misleading or incorrect answers.
@@
Although extensive research has focused on improving OCR accuracy and post-processing correction techniques, the specific impact of OCR noise on QA tasks remains underexplored. Previous studies have examined challenges related to OCR in information retrieval (IR) (de Oliveira et al., 2023), historical text processing (Piotrowski, 2012), and entity recognition (Grover et al., 2008). However, a systematic investigation of how OCR errors affect QA model performance is still lacking. Additionally, although large language models (LLMs) have been increasingly used to process OCR text, (Thomas et al., 2024; Madarász et al., 2024), their robustness in handling QA tasks with noisy OCR text remains an open research question.
@@
In this paper, we address the critical gap in understanding how large language models (LLMs) perform in question-answering (QA) tasks when dealing with noisy OCR-generated text. Specifically, we analyze the robustness of LLMs in QA tasks using RawOCR text and introduce MultiOCR-QA  [^2] The dataset is freely available at: https://github.com/DataScienceUIBK/MultiOCR-QA, a new multilingual QA dataset constructed from the ICDAR dataset, covering English, French, and German. This dataset includes both RawOCR text (OCR-generated text with errors) and CorrectedOCR text (ground truth), allowing a direct comparison of QA performance under different text quality conditions. To generate contextually relevant question-answer pairs from historical text excerpts, we leverage instruction-fine-tuned LLMs. We then systematically evaluate the impact of different types of OCR errors—insertions, deletions, and substitutions—on QA model performance, offering new insights into the strengths and limitations of LLMs when processing noisy historical data. Figure 1 illustrates an example from the MultiOCR-QA dataset, showcasing various OCR errors and their impact on text accuracy. It presents QA pairs and responses using both CorrectedOCR and RawOCR text as context, demonstrating the effects of OCR noise on QA performance.
@@
  • We introduce MultiOCR-QA, a new multilingual QA dataset from historical texts in English, French, and German, featuring both raw and corrected OCR text, allowing direct comparison under varying text quality conditions.
  • Using MultiOCR-QA, we then conduct a comprehensive evaluation of LLM robustness against noisy OCR text analyzing their impact on QA performance.
  • We also categorize different types and levels of OCR errors, evaluating their individual impact on QA performance, providing insights into how LLMs handle OCR-related challenges for different types of error.
Figure 2. The Pipeline for MultiOCR-QA generation: Arrows represent the output quantity based on the number of documents and callouts illustrate the statistics for English, French and German Language. Ps and QAs denote the number of paragraphs and question-answer pairs, respectively. Note, to supplement the data for English Language 10,000 questions from ChroniclingAmericaQA were included.
@@ 2. Related Work
Researchers have extensively studied the limitations of OCR for historical documents and its impact on information retrieval (IR). Croft et al. (1994) and Traub et al. (2015) examined how OCR errors reduce retrieval effectiveness. Chiron et al. (2017) found that 7% of the relevant documents were missed due to OCR misrecognition, demonstrating the risk of failure in matching noisy texts to user queries. While these studies highlight OCR challenges, they focus primarily on document retrieval and not on question answering (QA), which requires a more fine-grained understanding of text.
@@
Despite these insights into OCR’s effects on IR and NLP tasks, research on its impact on question answering remains limited. In the context of historical document collections, the only existing QA dataset, ChroniclingAmericaQA (Piryani et al., 2024), focuses primarily on creating a QA dataset from historical newspapers rather than systematically analyzing how different types of OCR errors affect QA performance. While studies on document retrieval and IR highlight OCR-related challenges, a comprehensive investigation into QA performance under different types and severity levels of OCR errors is still lacking. Our work fills this gap by introducing a multilingual QA dataset (MultiOCR-QA) and providing a detailed evaluation of large language models (LLMs) on the raw OCR text of MultiOCR-QA.
@@ 3.1. Data Collection
In this section, we describe the process of collecting documents to generate question-and-answer pairs to carry out our study. We utilized the ICDAR 2019 POST-OCR Text Correction dataset [^3] https://sites.google.com/view/icdar2019-postcorrectionocr (Rigaud et al., 2019) as our primary historical text source. We selected this dataset as it contains CorrectedOCR text (Ground Truth Text) along with RawOCR text. The dataset contains 22 million OCR-processed characters and corresponding ground truth text for ten European languages (English, French German, Finish, Spanish, Dutch, Czech, Bulgarian, Slovak, and Polish). In our study, we focused on English, French, and German.
Language Specific Data Collection: We extracted text in these three languages from the ICDAR 2019 dataset. The texts originally came from various historical document repositories.
@@
  • French: For the French language, the ICDAR 2019 dataset provides a collection from three sources: the HIMANIS [^4] https://www.himanis.org Project, IMPACT - National Library of France, and the RECEIPT [^5] http://findit.univ-lr.fr/ dataset. The original dataset contained 2,800 files.
@@
Furthermore, the ICDAR dataset, originally intended for post-OCR correction, contained special alignment symbols (e.g., @, #) to map the RawOCR text to its ground-truth counterpart. We removed the alignment symbols from the ground-truth text before generating questions. We also excluded files where the ground-truth text was missing, resulting in the removal of 16 files for English, 3 files for French, and none for German. This preprocessing step ensured that all QA pairs were generated from text that had both CorrectedOCR text and RawOCR text [^9] We will sometimes use CorrectedOCR and RawOCR to denote the text that was subject to post-OCR error correction and one without such correction, respectively..
@@ 3.2. Question-Answer Generation
For fine-tuning the model in English, we used the SQuAD v1 dataset (Rajpurkar et al., 2016). We randomly selected 2,067 paragraphs and 10,570 questions from the development set and 3,000 paragraphs and 13,894 questions from the test set. In total, we fine-tuned the LLaMA-3.1-70B instruct model for English QA dataset generation using 5,067 paragraphs and 24,464 questions. Similarly, for the French language, we fine-tuned the model on the FQuAD dataset (d’Hoffschmidt et al., 2020), utilizing both the validation and training sets, comprising 5,689 paragraphs and 23,919 questions. For the German language, we fine-tuned the model on the GermanQuAD dataset (Möller et al., 2021) using the training and validation sets, which consisted of 3,014 paragraphs and 13,722 questions. This fine-tuning step ensured that the model accurately generated factoid-style QA pairs, reducing instances of open-ended or conversational questions.
@@
English Answer Extraction Prompt System: You are an expert at extracting key information from text. Your goal is to identify spans of text that are likely to serve as answers to potential questions based on the input passage. Focus on meaningful, distinct, and diverse snippets such as entities, nouns, verbs, adjectives, numbers, dates, and phrases. Avoid redundancy and ensure the answers are diverse, representing key information in the passage. User: Given the passage below, extract several candidate spans that are likely to be answers to potential questions. Write only the extracted answers, separated by a semicolon (;). Passage : {context}
@@
English Question Generation Prompt System: You are an expert at generating standalone questions based on the provided passage. Your goal is to create a clear, relevant, and self-contained question that aligns with the information in the passage. The question should not explicitly reference the passage or require additional information to be understandable. Ensure the question is concise, well-structured, and meaningful. User: Based on the passage below, please generate a question that is relevant to the information provided. The question should be standalone, clear, and understandable without referencing the passage directly. The answer to the question should be [answer]. Passage : {context}
@@ 4. Dataset Analysis
After applying all the filtering steps, we obtained the final dataset, comprising 50,079 question-answer pairs. Due to the relatively low number of English-language documents in the ICDAR 2019 dataset, the initial MultiOCR-QA dataset contained only 875 QA pairs for English. To balance the distribution of QA pairs across languages, we incorporated 10,000 additional QA pairs from the ChroniclingAmericaQA dataset (Piryani et al., 2024). That dataset, sourced from the American historical newspaper collection called Chronicling America [^10] https://chroniclingamerica.loc.gov/about/, spans over 120 years from 1800-1920. It includes both ground-truth and RawOCR text, making it particularly relevant to our study. Consequently, the final MultiOCR-QA dataset consists of documents from historical newspapers, books, and centuries-old scripts, with 10,875 QA pairs in English, 10,004 in French, and 39,200 in German. The dataset statistics, including average paragraph length, question length, and answer length, are presented in Table 1.
@@
#QA pairs 10,875 10,004 39,200
#Paragraphs 6,525 1,670 9,075
Average CorrecteddOCR paragraph length (words) 219.09 297.53 212.86
Average RawOCR paragraph length (words) 233.25 335.73 193.23
Average question length (words) 10.98 8.73 8.08
Average answer length (words) 2.05 3.12 5.63
Average questions per paragraph 1.67 5.99 4.32
@@
4.1. Quantifying and Filtering OCR Noise

### v2

# Evaluating Robustness of LLMs in Question Answering on Multilingual Noisy OCR Data

@@ Abstract.
Optical Character Recognition (OCR) plays a crucial role in digitizing historical and multilingual documents, yet OCR errors - imperfect extraction of text, including character insertion, deletion, and substitution can significantly impact downstream tasks like question-answering (QA). In this work, we conduct a comprehensive analysis of how OCR-induced noise affects the performance of Multilingual QA Systems. To support this analysis, we introduce a multilingual QA dataset MultiOCR-QA, comprising 50K question-answer pairs across three languages, English, French, and German. The dataset is curated from OCR-ed historical documents, which include different levels and types of OCR noise. We then evaluate how different state-of-the-art Large Language models (LLMs) perform under different error conditions, focusing on three major OCR error types. Our findings show that QA systems are highly prone to OCR-induced errors and perform poorly on noisy OCR text. By comparing model performance on clean versus noisy texts, we provide insights into the limitations of current approaches and emphasize the need for more noise-resilient QA systems in historical digitization contexts.
@@
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXconference: The 34th ACM International Conference on Information and knowledge Management; November 10–14, 2025; Seoul, Koreaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Question answeringccs: Information systems Content analysis and feature selection
@@ 1. Introduction
Optical Character Recognition (OCR) technology has played a crucial role in digitizing and providing access to historical texts. Over the past decade, significant advancements in OCR have improved text recognition accuracy, leading to the development of large-scale digital libraries of historical texts (TERRAS, 2011). These libraries serve as valuable resources for researchers, professionals and the general public, enabling access to old manuscripts, newspapers, and other archival materials. Historical documents hold a wealth of knowledge, offering insights into past events, cultures, and people. Many professionals such as historians, journalists, or sociologists rely on these heritage collections for various research and analysis tasks.
Historical corpora have been already used as an underlying text for numerous natural language processing (NLP) tasks, including Named Entity Recognition (NER), topic modeling, text classification, neural ranking, and understanding semantic changes in language over time (Ehrmann et al., 2020, 2023; Yang et al., 2011; Liebeskind and Liebeskind, 2020; Subramani et al., 2020). While these tasks offer valuable insights, it is expected that automatic question answering (QA) systems become an important interface to the vast historical collections. QA systems enable direct and intuitive access to specific information, offering a powerful means of analyzing, searching and understanding historical texts. By retrieving precise answers to user queries, QA can significantly enhance the accessibility, interpretability, and utility of historical corpora, making them more easily accessible and more useful for scholars, researchers, and the general public alike.
@@
Despite the advancements in OCR technology, significant challenges such as character misrecognition and structural inconsistencies persist. OCR-generated text, referred as RawOCR text in this paper, often contains errors due to the degraded conditions and non-standard character of historical documents. Factors like faded ink and paper, irregular fonts, physical damage, and printing inconsistencies cause recognition errors that negatively impact downstream NLP tasks such as information retrieval, machine translation, and QA systems. Since QA models heavily depend on the quality of the input text, errors in raw OCR text can significantly affect the accuracy and reliability of generated answers.
@@
For instance, in German, a passage with OCR error ”Der Bericht der Lagsatzungsgesandtschaft wird verlesen undvon Hrn. Bürgermeistet Mousson als erstem Gesandten desStandes Zürich mit einigen Bemerkungen begleitet. [^1] English Translation: The report of the legislative mission is read out and accompanied by a few comments from Mr. Mayor Mousson, the first envoy of the Zurich state.” contains multiple insertion, deletion, and substitution errors, such as ”Lagsatzungsgesandtschaft” (should be ”Tagsatzungsgesandtschaft”) and ”Bürgermeistet” (should be ”Bürgermeister”). When a QA model encounters such a noisy OCR text, it may generate an incorrect answer. For example, given the question: ”Wer hat den Bericht der Tagsatzungsgesandtschaft verlesen?” (Who read the report of the parliamentary delegation?) the model incorrectly responds ”Der Bericht der Tagsatzungsgesandtschaft wurde von Hrn. Bürgermeistet Mousson verlesen.” (The report of the parliamentary delegation was read out by Mr. Mayor Mousson.) which retains the OCR error and potentially misleads the QA system. This example highlights how even minor OCR errors can significantly affect QA accuracy, resulting in misleading or incorrect answers.
@@
Although extensive research has focused on improving OCR accuracy and post-processing correction techniques, the specific impact of OCR noise on QA remains largely unexplored. Previous studies have examined challenges related to OCR in information retrieval (IR) (de Oliveira et al., 2023), historical text processing (Piotrowski, 2012), and entity recognition (Grover et al., 2008). However, a systematic investigation of how OCR errors affect QA model performance is still missing. Furthermore, our study contributes also to the body of work on robustness of LLMs against noisy inputs such as noisy prompts (Wang et al., 2023), however we do it in a quite novel yet realistic setting.
@@
In this paper, we address this gap in understanding how large language models (LLMs) perform in question-answering (QA) tasks when dealing with noisy OCR-generated text. In particular, we first introduce MultiOCR-QA  [^2] The dataset and code will be made available after paper publication. a new multilingual QA dataset covering English, French, and German. This dataset includes both RawOCR text (OCR-generated text with errors) and CorrectedOCR text (ground truth), allowing for a direct comparison of QA performance under different text quality conditions. To generate contextually relevant question-answer pairs from historical text excerpts, we leverage instruction-fine-tuned LLMs. We then systematically evaluate the impact of different types of OCR errors—insertions, deletions, and substitutions—on QA model performance, offering new insights into the strengths and limitations of LLMs when processing noisy historical data. Figure 1 illustrates an example from the MultiOCR-QA dataset with various OCR errors impacting text accuracy. It presents QA pairs and the corresponding responses for questions when using CorrectedOCR and RawOCR text as input context to Gemma 2–27b, demonstrating the effects of OCR noise on QA performance. For instance, Question 1: ”What did the man say happened to his leg?” produces different answers depending on whether clean or noisy text was input. When using the CorrectedOCR context, the model produces the response: ”He said his sense of the leg was gone.” In contrast, the response based on the RawOCR context is: ”The fence of his leg is gone,” demonstrating how OCR-induced noise can significantly alter the meaning of the answer and degrade the QA accuracy.
@@
  • We introduce MultiOCR-QA, a new multilingual QA dataset from historical texts in English, French, and German, featuring both raw and corrected OCR text, that allows direct comparison under varying text quality conditions.
  • Using MultiOCR-QA, we conduct a comprehensive evaluation of LLM robustness against noisy OCR text analyzing and quantifying their impact on QA performance.
  • We categorize different types and degrees of OCR errors for the purpose of evaluating their individual impact on QA performance. This allows providing comprehensive insights into how LLMs handle OCR-related challenges for different types and severity of errors in each of the studied languages.
Figure 2. The Pipeline for MultiOCR-QA generation: Arrows represent the output quantity based on the number of documents and callouts illustrate the statistics for English, French and German Language. Ps and QAs denote the number of paragraphs and question-answer pairs, respectively.
@@ 2. Related Work
Several research works have been carried to study the limitations of OCR for historical documents and its impact on information retrieval (IR). Croft et al. (1994) and Traub et al. (2015) examined how OCR errors reduce retrieval effectiveness. Chiron et al. (2017) found that 7% of the relevant documents were missed due to OCR misrecognition, demonstrating the risk of failure in matching noisy texts to user queries. While these studies highlight OCR challenges, they focus primarily on document retrieval and not on question answering (QA), which requires a more fine-grained understanding of text.
@@
Despite these insights into OCR’s effects on IR and NLP tasks, research on its impact on question answering remains limited. In the context of historical document collections, the only existing QA dataset, ChroniclingAmericaQA (Piryani et al., 2024), focuses primarily on creating a QA dataset from historical newspapers rather than systematically analyzing how different types of OCR errors affect QA performance. While studies on document retrieval and IR highlight OCR-related challenges, a comprehensive investigation into QA performance under different types and severity levels of OCR errors is still missing. Our work fills this gap by introducing a multilingual QA dataset (MultiOCR-QA) and providing a detailed evaluation of large language models (LLMs) on the raw OCR text of MultiOCR-QA.
@@ 3.1. Data Collection
We first describe the process of collecting documents to generate question-answer pairs for our study. Although several historical text datasets exist, such as the IMPACT Project  [^3] https://www.digitisation.eu/impact-dataset/ and the Europeana Newspaper Project  [^4] http://www.europeana-newspapers.eu/, which include digitized images of historical documents alongside their OCR-processed text, these lack systematically aligned ground truth corrections for a robust analysis of OCR noise effects. The ICDAR 2019 POST-OCR Text Correction dataset [^5] https://sites.google.com/view/icdar2019-postcorrectionocr (Rigaud et al., 2019) is an exception here as it provides both RawOCR text along with its aligned ground truth (called CorrectedOCR), making it especially suitable for our research objectives. Furthermore, several English, French, and German document sources included in the IMPACT and Europeana projects are already incorporated into the ICDAR 2019 dataset, offering a diverse collection of cleaned resources for historical OCR-based studies.
ICDAR 2019 dataset includes over 22 million OCR-processed characters and their corresponding, aligned ground truth across several European languages. We focus on English, French, and German languages due to several reasons. First, French and German are among the most well-represented languages in the ICDAR 2019 dataset regarding their document numbers. Second, all three languages are high-resource and well-supported by the current large language models (LLMs), which increases the reliability of QA generation and evaluation (Li et al., 2025). Since noisy OCR text is expected to pose already a significant challenge for LLMs, incorporating low-resource languages, where LLM performance tends to be generally weaker, might introduce additional complexity. Third, each of these languages has an existing, publicly available QA dataset like SQuAD (Rajpurkar et al., 2016) for English, FQuAD (d’Hoffschmidt et al., 2020) for French, and GermanQuAD (Möller et al., 2021) for German, which enable effective instruction fine-tuning of LLMs for high-quality QA pair generation. Therefore, we limited our study to these three languages, intending to expand to low-resource and typologically diverse languages in the future work.
Language Specific Data Collection: The texts in ICDAR 2019 dataset originally came from various historical document repositories.
@@
  • French: For the French language, the ICDAR 2019 dataset provides a collection obtained from three sources: the HIMANIS [^6] https://www.himanis.org Project, IMPACT - National Library of France, and the RECEIPT [^7] http://findit.univ-lr.fr/ dataset. The original dataset contained 2,800 files.
@@
Furthermore, the ICDAR dataset, originally intended for post-OCR correction, contained special alignment symbols (e.g., @, #) to map the RawOCR text to its ground-truth counterpart. We removed these symbols from the ground-truth text before generating questions. We also excluded files where the ground-truth text was missing, resulting in the removal of 16 files for English, 3 files for French, and none for German. This preprocessing step ensured that all QA pairs were generated from text that had both CorrectedOCR text and RawOCR text.
@@ 3.2. Question-Answer Generation
For fine-tuning the model in English, we used the SQuAD v1 dataset (Rajpurkar et al., 2016). We randomly selected 2,067 paragraphs and 10,570 questions from the development set and 3,000 paragraphs and 13,894 questions from the test set. Similarly, for the French language, we fine-tuned the model on the FQuAD dataset (d’Hoffschmidt et al., 2020), utilizing both the validation and training sets, comprising 5,689 paragraphs and 23,919 questions. For the German language, we fine-tuned the model on the GermanQuAD dataset (Möller et al., 2021) using the training and validation sets, which consisted of 3,014 paragraphs and 13,722 questions. This fine-tuning step ensured that the model accurately generated factoid-style QA pairs, reducing instances of open-ended or conversational questions.
@@
English Answer Extraction Prompt System Prompt: You are an expert at extracting key information from text. Your goal is to identify spans of text that are likely to serve as answers to potential questions based on the input passage. Focus on meaningful, distinct, and diverse snippets such as entities, nouns, verbs, adjectives, numbers, dates, and phrases. Avoid redundancy and ensure the answers are diverse, representing key information in the passage. User Prompt: Given the passage below, extract several candidate spans that are likely to be answers to potential questions. Write only the extracted answers, separated by a semicolon (;). Passage : {context}
@@
English Question Generation Prompt System Prompt: You are an expert at generating standalone questions based on the provided passage. Your goal is to create a clear, relevant, and self-contained question that aligns with the information in the passage. The question should not explicitly reference the passage or require additional information to be understandable. Ensure the question is concise, well-structured, and meaningful. User Prompt: Based on the passage below, please generate a question that is relevant to the information provided. The question should be standalone, clear, and understandable without referencing the passage directly. The answer to the question should be [answer]. Passage : {context}
@@ 4. Dataset Analysis
After applying all the filtering steps, we obtained the final dataset, comprising 50,079 question-answer pairs.
The dataset statistics, including average paragraph length, question length, and answer length, are presented in Table 1.
@@
#QA pairs 875 10,004 39,200
#Paragraphs 123 1,670 9,075
Avg. CorrectedOCR paragraph length (words) 271.73 297.53 212.86
Avg. RawOCR paragraph length (words) 263.46 335.73 193.23
Avg. question length (words) 8.60 8.73 8.08
Avg. answer length (words) 2.05 3.12 5.63
Avg. questions per paragraph 7.11 5.99 4.32
@@
4.1. Quantifying and Filtering Noise